OpenStudy (anonymous):

1/(6t+1)^1/2 dt evaluated from 0 to 4

OpenStudy (anonymous):

The integral of that you mean?

OpenStudy (anonymous):

Yes, sorry I wasn't clear

OpenStudy (anonymous):

Definite Integral

myininaya (myininaya):

let u=6t+1

OpenStudy (anonymous):

Use a u substation as myininaya has just suggested ;p

OpenStudy (anonymous):

substitution I mean.

OpenStudy (anonymous):

I have to use the Fundamental Theorem of Calculus

myininaya (myininaya):

we will

OpenStudy (anonymous):

Ok

myininaya (myininaya):

there is no other way unless you used riemann sums but lets not mention him anymore

OpenStudy (anonymous):

haha

myininaya (myininaya):

using riemann sums can also be very difficult to evaluate the limit

myininaya (myininaya):

or maybe impossible

OpenStudy (anonymous):

Well I'm teaching myself Integrals so I'm not quite that far yet.

OpenStudy (anonymous):

I'm just having trouble finding where to start with these functions.

myininaya (myininaya):

do you know substation?

OpenStudy (anonymous):

substitution ;p

myininaya (myininaya):

stupid thingy i hit whatever popped up because i hate spelling substitution

OpenStudy (anonymous):

You have to learn u substitutions to do these they allow you to undo the chain rule.

OpenStudy (anonymous):

Ok, So after the substitution, What's my next step?

myininaya (myininaya):

if u=6t+1 then du/dt=.....

OpenStudy (anonymous):

6

myininaya (myininaya):

right do du=....

OpenStudy (anonymous):

6 dt?

myininaya (myininaya):

right

OpenStudy (anonymous):

Then I hit both sides with integrals, correct?

OpenStudy (anonymous):

No. Then you substitute 6t +1 as u, and your dt as 1/6 du.

OpenStudy (anonymous):

And take the integral with respect to u (though you also have to change your limits)

OpenStudy (anonymous):

Oh, ok, that's where I get lost.

myininaya (myininaya):

polpak divided both sides by 6 of that equation du=6dt

OpenStudy (anonymous):

Well 6t + 1 = u the 1/(6t+1) = 1/u

myininaya (myininaya):

you dont have to change the limits you can just leave it blank and come back after finding the antiderivative in terms of x in using the limits that we started with

OpenStudy (anonymous):

and dt = 1/6 du so \[\int_0^4 \frac{1}{6t+1}dt = \int_1^{25} \frac{1}{u}(\frac{1}{6} du)\]

OpenStudy (anonymous):

Since 1/6 is just a multiplicative constant you can pull it out front.

myininaya (myininaya):

when i say x i meant t lol

OpenStudy (anonymous):

Hmmmm...

OpenStudy (anonymous):

Ok, let's try this one step at a time. Let u = 6t +1 \[\implies du = 6dt \implies dt = \frac{1}{6}du\] \[t = 0 \implies u = 0 + 1 = 1\] \[t = 4 \implies u = 6(4) + 1 = 25\] Therefore \[\int_0^4 \frac{1}{6t+1} dt = \int_1^{25} \frac{1}{u}(\frac{1}{6}du) \] \[= \frac{1}{6} \int_1^{25} \frac{1}{u}du\]

OpenStudy (anonymous):

Make more sense?

OpenStudy (anonymous):

The only thing I'm confused about is where \[\int\limits_{1}^{25}\] comes from

OpenStudy (anonymous):

See the t=0 and t=4 lines? If we are evaluating t from 0 to 4, that means u goes from 1 to 25.

OpenStudy (anonymous):

Ohhh, okay, I see

OpenStudy (anonymous):

But, the original equation it's \[\sqrt{6t+1}\] What does that change in your answer?

myininaya (myininaya):

i just posted it

myininaya (myininaya):

1 over that right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Wow, thanks for that effort in explaining, that helps so much!

OpenStudy (anonymous):

Oh, sorry didn't see the exponent. It changes things some in that it's \(1/u^2\) instead of 1/u, but that just changes the end result, not too much in what I posted.

OpenStudy (anonymous):

Ok, thank you both, very much appreciated guys, I have another question as well if you want to answer it.

OpenStudy (anonymous):

The population of a certain community (t) years after the year 2000 is given by \[P(t)= e ^{0.2t}\div4+e ^{0.2t} million people\] what was the average population from 2000-2010?

myininaya (myininaya):

for 2000, t=0 is that right? t=1?

OpenStudy (anonymous):

I'm not sure, that's all I was given and I'm not too sharp in this subject of integrals yet, sorry.

myininaya (myininaya):

polpak come back? lol

OpenStudy (anonymous):

Haha

myininaya (myininaya):

we can work on integrating P(t) together let me look at it, and I will scan it in k?

OpenStudy (anonymous):

Thank you so much! Please! lol

myininaya (myininaya):

hey so we have [e^(0.2t)]/4 + e^(0.2t) ={5e^(0.2t)}/4?

OpenStudy (anonymous):

I'm so stressed out. Lol

OpenStudy (anonymous):

Where did the 5 come from?

myininaya (myininaya):

which 5?

OpenStudy (anonymous):

Oh duh, I was looking at it too hard.

myininaya (myininaya):

so you are teaching yourself?

OpenStudy (anonymous):

Yes

myininaya (myininaya):

thats cool. we need to figure out what limits to use? its ether 1 to 11 or 0 to 10 depending if we start 2000 with t=1 or t=0 i will get polpak and see if he knows

OpenStudy (anonymous):

Ok, so how do you determine what limits to use?

myininaya (myininaya):

well it wants to know the average population from 2000 to 2010 so from there we can figure out the limits

OpenStudy (anonymous):

Ok, so the t=1 or t=0 would represent the year 2000, as the 10/11 would represent the 10 year time frame from 2000.

myininaya (myininaya):

right

OpenStudy (anonymous):

Alrighty, so what difference would it make to use either one of the limits?

OpenStudy (anonymous):

Sorry , what's up?

OpenStudy (anonymous):

Oh, 2000 is t=0

myininaya (myininaya):

ok cool lol

OpenStudy (anonymous):

"(t) years after the year 2000" so if t = 0 the year is 2000.

OpenStudy (anonymous):

t=10 would be 2010

myininaya (myininaya):

ok so 2010 would be t=10 so we have our upper and lower limit just use the fundamental thm of calculus on the antiderivative we found

myininaya (myininaya):

right i was not sure at first if it was 1 to 11 or 0 to 10 thanks polpak

OpenStudy (anonymous):

np =)

OpenStudy (anonymous):

Thank you polpak!

myininaya (myininaya):

so you do [25e^(0.2*10)]/4 - [25e^(0*10)]/4

OpenStudy (anonymous):

What did you get after evaluation?

OpenStudy (anonymous):

I got 40 million people